OpenStudy (anonymous):

$720, 4.25%, 3 months. find the interest rate

OpenStudy (anonymous):

can i have an answer

OpenStudy (anonymous):

4.25% is 0.0425 in decimal....multiply by $720, and interest rate is $30.60

OpenStudy (anonymous):

how is that possible if you started with 720 dollars and ended up with 30 dollars

OpenStudy (anonymous):

Hold on, the interest rate formula is as so: A = P(1+rt), where A is interest rate, P is your starting amount or principal, r is the rate as a decimal, and t is time. 3 months divided by 12 months per year is 1/4, or 0.25. So plug in everything: A = 720(1+0.0425 * 0.25) = $727.65, however your interest rate should simply be 4.25%

OpenStudy (anonymous):

This is used to calculate the principal plus interest